### v4

@@ Abstract
Recent advances in wave modeling use sufficiently accurate fine solver outputs to train a neural network that enhances the accuracy of a fast but inaccurate coarse solver. In this paper we build upon the work of Nguyen and Tsai (2023) and present a novel unified system that integrates a numerical solver with a deep learning component into an end-to-end framework. In the proposed setting, we investigate refinements to the network architecture and data generation algorithm. A stable and fast solver further allows the use of Parareal, a parallel-in-time algorithm to correct high-frequency wave components. Our results show that the cohesive structure improves performance without sacrificing speed, and demonstrate the importance of temporal dynamics, as well as Parareal, for accurate wave propagation.
@@ 1 Introduction
Simulating wave propagation is an important task for many applications. For instance, when searching natural gas, wave propagation computations form the forward part of a numerical method for solving the inverse problem of geophysical inversion. By accurately analyzing the reflections and transmissions generated by complex media discontinuities, it becomes possible to characterize underground formations. However, traditional numerical computations often demand a computationally expensive fine grid representation to guarantee stability.
@@ 2 Finite-Difference-Based Wave Propagators
We consider absorbing or periodic boundary conditions and smooth initial conditions. Following the setup in nguyen2022numerical , let $Q_{h}u$ denote a numerical approximation of $\Delta u$ with discretized spatial and temporal domains, i.e.,
@@
Figure 1: Detailed schematic of E2E-JNet3 (adapted from nguyen2022numerical ). Each convolutional block (blue) encompasses a 3x3 convolutional layer ($\text{groups}=3$, $\text{padding}=1$), followed by a batch normalization and a ReLU activation function. To reduce the resolution, the 3rd and 6th convolutional block (red) employs a stride of two. Connectivity within the network is depicted by arrows, with the dashed arrow specifically indicating a single application of $\mathcal{G}_{\Delta t^{\star}}$.
@@
A simple model with bilinear interpolation (E2E-V, $\mathcal{I}^{0}\Psi_{\Delta t^{\star}}^{\theta}\mathcal{R}$) is used as a baseline. Each variant changes the baseline by exactly one aspect. This allows us to isolate the effect of each modification on the performance of the architecture.
The four investigated end-to-end models $\Psi_{\Delta t^{\star}}^{\theta}\coloneqq\mathcal{I}_{\Delta t^{\star}}^{~{}\theta}\mathcal{G}_{\Delta t^{\star}}\mathcal{R}$ are:
@@ 3 Data Generation Approaches
    An initial wave field $\mathfrak{u}_{0}=(u_{0},\partial_{t}u_{0})\in\mathfrak{F}_{u_{0}}$ is sampled from a Gaussian pulse,
    $\mathfrak{u}_{0}(x_{1},x_{2})=(u_{0},p_{0})=(e^{-\frac{(x+\tau)^{2}}{\sigma^{2}}},~{}0),$ (3)
    with $x\in[-1,1]^{2}$ and $\frac{1}{\sigma^{2}}\sim\mathcal{N}(250,10)$. $\tau\in[-0.5,0.5]^{2}$ represents the displacement of the Gaussian pulse’s location from the center.
@@
    Every $\mathfrak{u}_{0}\in\mathfrak{F}_{u_{0}}$ is then propagated eight time steps $\Delta t^{\star}=0.06$ by $\mathcal{F}_{\Delta t^{\star}}$. We adopt the fine grid settings of the spatial ($\delta x=\frac{2}{128}$) and temporal resolution ($\delta t=\frac{1}{1280}$) from nguyen2022numerical .
@@
Wave Speeds.
We sample $c\in\mathfrak{F}_{c}$ from randomly chosen subregions of two synthetic geological structures, Marmousi marmousi and BP bp , that are mapped onto the spatial grid $h\mathbb{Z}^{2}\cap[-1,1]^{2}$. For testing, four manually modified media are added to examine rapid variations in wave speed:
  • Diagonal Ray: $c(x_{1},x_{2})=3-1.5~{}[|x_{1}+x_{2}|>0.3]$
  • Three Layers: $c(x_{1},x_{2})=2.5-0.7~{}[|x_{1}+x_{2}|>-0.4]*2~{}[|x_{1}+x_{2}|>-0.6]$
  • Wave Guide: $c(x_{1},x_{2})=3-0.9~{}\cos(\pi x_{1})$
[…]
Marmousi and BP profiles are drawn with a probability of $30\%$ each, and the other velocity profiles with a probability of $10\%$ each, respectively.
Dataset Split.
Different sets of hyperparameters are used to train on the training set and evaluate on the validation set. A total of $24$ grid searches are performed for three trials tuning the learning rate $\thicksim\{10^{-3},10^{-4}\}$, weight decay $\thicksim\{10^{-2},10^{-3}\}$, and batch size $\thicksim\{2^{6},2^{8}\}$. During validation and testing, the variants are iteratively applied to an initial condition sampled from $\mathcal{D}$ for eight steps. To ensure equal conditions, the training set is limited to $40{,}000$, and the validation and test set to $5{,}000$ data points.
@@ 3.1 Multi-Step Training
For each initial condition $\mathfrak{u}_{0}$, $\mathcal{F}_{\Delta t}$ is applied $N$ times with solutions denoted as $\mathfrak{u}_{n},\forall n\in\mathcal{U}_{1}\coloneqq\{0,\ldots,N\}$. In random order, $\Psi_{\Delta t}^{\theta}$ is applied to every $\mathfrak{u}_{n}$ for a random amount of steps $k\in\mathcal{U}_{2}\coloneqq\{1,\ldots,N-n\}$. Formally, the optimization problem can therefore be described as:
$\min_{\theta\in\mathbb{R}^{m\times n}}\mathcal{L}(\Psi_{\Delta t}^{\theta};\mathcal{D})=\min_{\theta\in\mathbb{R}^{m\times n}}\frac{1}{|\mathcal{D}|}\sum\limits_{\mathfrak{u}_{0}}\sum\limits_{n\sim\mathcal{U}_{1}\setminus{\{N\}}}\sum\limits_{\begin{subarray}{c}k\sim\mathcal{U}_{2}\\ n<k\leq N-n\end{subarray}}\lVert(\Psi_{\Delta t}^{\theta})^{k}\mathfrak{u}_{n}-(\mathcal{F}_{\Delta t})^{k}\mathfrak{u}_{n}\rVert_{E_{h}}^{2}.$ (6)
The norm $\lVert\cdot\rVert_{E_{h}}^{2}$ is the discretized energy semi-norm MSE as detailed in nguyen2022numerical . We draw both $n$ and $k$ from the uniform random distributions $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$, respectively. The novel dataset is denoted as $\mathcal{D}^{m}=\{(\nabla u_{n}^{k},c^{-2}(u_{n}^{k})_{t},c;\nabla u_{n+1},c^{-2}(u_{n+1})_{t})\}$.
@@
Since in the model’s initial, untrained phase, feature variations can be extreme and may lead to imprecise gradient estimations, we aim to accelerate convergence by weighting individual losses. Therefore, rather than drawing $k\in\mathcal{U}_{2}$ from a uniform distribution, we select values for $k$ according to a truncated standard normal distribution TN$(\mu,\sigma,a,b)$ from the sample space represented as $-\infty<a<b<\infty$. This focuses on minimizing the impact of errors in the early training stage. After every third epoch, the mean $\mu$ is increased by one to account for long-term dependencies. We refer to this dataset as $\mathcal{D}^{w,m}$.
@@ 5 Discussion
Figure 3: Comparing the NE2E-JNet3 model and three end-to-end JNet variants that differ in their training algorithms. Initial conditions and velocity profiles are sampled from $\mathcal{D}$ and the relative energy MSE results of 10 runs are averaged. As expected, all models show a bounded growth as the waves vanish due to absorbing boundary conditions.
@@ 6 Conclusion
All end-to-end variants, including the variants with training modifications, outperformed the modular framework of nguyen2022numerical . In particular, the lightweight end-to-end 3-level JNet (E2E-JNet3) performed reasonably well given its low computation cost, and was further improved through a weighted, multi-step training scheme ($\mathcal{D}^{w,m}$) to feature time-dependent wave dynamics without adding complexity to the model or substantially extending the training duration. Similarly, the Parareal iterations using the neural propagator trained by the “Parareal”-based data showed significant performance improvements over E2E-JNet3 without extensive additional computational cost due to parallelization.

### v5

@@ Abstract
In a variety of scientific and engineering domains, the need for high-fidelity and efficient solutions for high-frequency wave propagation holds great significance. Recent advances in wave modeling use sufficiently accurate fine solver outputs to train a neural network that enhances the accuracy of a fast but inaccurate coarse solver. In this paper we build upon the work of Nguyen and Tsai (2023) and present a novel unified system that integrates a numerical solver with a deep learning component into an end-to-end framework. In the proposed setting, we investigate refinements to the network architecture and data generation algorithm. A stable and fast solver further allows the use of Parareal, a parallel-in-time algorithm to correct high-frequency wave components. Our results show that the cohesive structure improves performance without sacrificing speed, and demonstrate the importance of temporal dynamics, as well as Parareal, for accurate wave propagation.
@@ 1 Introduction
Wave propagation computations form the forward part of a numerical method for solving the inverse problem of geophysical inversion. This involves solving the wave equation with highly varying sound speed many times in a most efficient way. For instance, by accurately analyzing the reflections and transmissions generated by complex media discontinuities, it becomes possible to characterize underground formations when searching for natural gas. However, traditional numerical computations often demand a computationally expensive fine grid to guarantee stability.
@@ 2 Finite-Difference-Based Wave Propagators
We consider smooth initial conditions and absorbing or periodic boundary conditions that lead to well-posed initial boundary value problems. Since we are interested in seismic exploration applications, both boundary conditions can be used to simulate the propagation of wave fields with initial energy distributed inside a compact domain. Following the setup in nguyen2022numerical , let $Q_{h}u$ denote a numerical approximation of $\Delta u$ with discretized spatial and temporal domains, i.e.,
@@
Figure 1: Detailed schematic of E2E-JNet3 (adapted from nguyen2022numerical ). Each convolutional block (blue) encompasses a 3x3 convolutional layer ($\text{groups}=3$, $\text{padding}=1$), followed by a batch normalization and a ReLU activation function. Connectivity within the network is depicted by arrows, with the dashed arrow specifically indicating a single application of $\mathcal{G}_{\Delta t^{\star}}$.
@@
A simple model with bilinear interpolation (E2E-V, $\mathcal{I}^{0}\Psi_{\Delta t^{\star}}^{\theta}\mathcal{R}$) is used as a baseline. Each variant changes the baseline by exactly one aspect. This allows us to isolate the effect of each architecture modification. The four investigated end-to-end models $\Psi_{\Delta t^{\star}}^{\theta}\coloneqq\mathcal{I}_{\Delta t^{\star}}^{~{}\theta}\mathcal{G}_{\Delta t^{\star}}\mathcal{R}$ are:
@@ 3 Data Generation Approaches
    An initial wave field $\mathfrak{u}_{0}=(u_{0},p_{0})\in\mathfrak{F}_{u_{0}}$ is sampled from a Gaussian pulse,
    $u_{0}=e^{-\frac{|x+\tau|^{2}}{\sigma^{2}}},p_{0}\equiv\partial_{t}u_{0}=0$ (3)
    with $x\in[-1,1]^{2}$, $\frac{1}{\sigma^{2}}\sim\mathcal{N}(250,10)$ and the initial velocity field $p_{0}$. $\tau\in[-0.5,0.5]^{2}$ is the displacement of the Gaussian pulse’s location from the center.
@@
    Every $\mathfrak{u}_{0}\in\mathfrak{F}_{u_{0}}$ is then propagated eight time steps $\Delta t^{\star}=0.06$ by $\mathcal{F}_{\Delta t^{\star}}$. We adopt the fine grid settings for the spatial ($\delta x=\frac{2}{128}$) and temporal resolution ($\delta t=\frac{1}{1280}$) from nguyen2022numerical .
@@
Wave Speeds
$c\sim\mathfrak{F}_{c}$ are sampled from randomly chosen subregions of two synthetic geological structures, Marmousi marmousi and BP bp , that are mapped onto the spatial grid $h\mathbb{Z}^{2}\cap[-1,1]^{2}$ (see Figure 2). Four manually modified media (cf. nguyen2022numerical ) are added during testing to examine rapid variations in wave speed.
Figure 2: Velocity profiles. Brighter colors indicate higher velocity, while randomly chosen subregions are shown in red squares. Marmousi and BP profiles are drawn with a probability of $30\%$ each, and the other velocity profiles with a probability of $10\%$ each, respectively.
@@ 3.1 Multi-Step Training
For each initial condition $\mathfrak{u}_{0}$, $\mathcal{F}_{\Delta t}$ is applied $N$ times with solutions denoted as $\mathfrak{u}_{n},\forall n\in\mathcal{U}_{1}\coloneqq\{0,\ldots,N\}$. In random order, $\Psi_{\Delta t}^{\theta}$ is applied to every $\mathfrak{u}_{n}$ for a random amount of steps $k\in\mathcal{U}_{2}\coloneqq\{n+1,\ldots,N-n\}$. Formally, the optimization problem can therefore be described as:
$\min_{\theta\in\mathbb{R}^{m\times n}}\mathcal{L}(\Psi_{\Delta t}^{\theta};\mathcal{D})=\min_{\theta\in\mathbb{R}^{m\times n}}\frac{1}{|\mathcal{D}|}\sum\limits_{\mathfrak{u}_{0},c}\sum\limits_{n\sim\mathcal{U}_{1}\setminus{\{N\}}}\sum\limits_{\begin{subarray}{c}k\sim\mathcal{U}_{2}\\ n<k\leq N-n\end{subarray}}\lVert(\Psi_{\Delta t}^{\theta})^{k}\mathfrak{u}_{n}-(\mathcal{F}_{\Delta t})^{k}\mathfrak{u}_{n}\rVert_{E_{h}}^{2}.$ (6)
The norm $\lVert\cdot\rVert_{E_{h}}^{2}$ is the discretized energy semi-norm MSE as detailed in nguyen2022numerical . We draw both $n$ and $k$ from the uniform random distributions, i.e., $n\sim\mathcal{U}_{1}$ and $k\sim\mathcal{U}_{2}$, respectively. The novel dataset is denoted as $\mathcal{D}^{m}=\{(\nabla u_{n}^{k},c^{-2}(u_{n}^{k})_{t},c;\nabla u_{n+1},c^{-2}(u_{n+1})_{t})\}$.
@@
Since in the model’s initial, untrained phase, feature variations can be extreme and may lead to imprecise gradient estimations, we aim to accelerate convergence by weighting individual losses. Therefore, rather than drawing $k\sim\mathcal{U}_{2}$ from a uniform distribution, we select values according to a truncated normal distribution TN$(\mu,\sigma,a,b)$ from the sample space represented as $-\infty<a<b<\infty$. This focuses on minimizing the impact of errors in the early training stage. After every third epoch, the mean $\mu$ is increased by one to account for long-term dependencies. We refer to this dataset as $\mathcal{D}^{w,m}$.
@@ 5 Discussion
Figure 4: Comparing the NE2E-JNet3 model and three end-to-end JNet3 variants that differ in their training algorithms. Initial conditions and velocity profiles are sampled from $\mathcal{D}$ and the relative energy MSE results of 10 runs are averaged. As expected, all models show a bounded growth as the waves vanish due to absorbing boundary conditions.
@@ 6 Conclusion
All end-to-end variants, including the variants with training modifications, outperformed the modular framework of nguyen2022numerical . In particular, the lightweight end-to-end 3-level JNet (E2E-JNet3) performed reasonably well given its low computation cost, and was further improved through a weighted, multi-step training scheme ($\mathcal{D}^{w,m}$) to feature time-dependent wave dynamics without adding complexity to the model or substantially extending the training duration. Similarly, the Parareal iterations using the neural propagator trained by the Parareal-based data showed significant performance improvements over E2E-JNet3 without extensive additional computational cost due to parallelization.